OpenStudy (anonymous):

algebra help please? just a few questions im stuck on out of 48

OpenStudy (anonymous):

12. Solve the equation. 3x – 10 = 3(x – 4) + 2 x = 8 x = 0 x = infinitely many solutions x = no solution

OpenStudy (anonymous):

x=no solution

OpenStudy (anonymous):

Do you mind helping with a few more ? 22. Solve. l x - 1/5 l = 3 x = – 14/5and x = 16/5 x =14/5 and x = – 16/5 x = – 14/5 and x = – 16/5 x = 14/5 and x = 16/5

OpenStudy (anonymous):

@khuangg its preferred if you help the asker through the problem and not just post the solution.

OpenStudy (anonymous):

Dr.Pepper do you know how to separate an absolute value?

OpenStudy (anonymous):

I am sorry but for whatever reason if you replied to me, which it told me you did, then it is not appearing on my screen could you please repost it?

OpenStudy (anonymous):

oo sorry i said yes kind of but i get confused

OpenStudy (anonymous):

okay can you show me what the LHS will be when you separate it?

OpenStudy (anonymous):

If you need help with this part tell me but I prefer to let you try first and help you when you stumble, if you are stumbling post that and I will work through it with you

OpenStudy (anonymous):

sorry i was looking through my book but i cant find it >_<

OpenStudy (anonymous):

In general an absolute value means that if the function inside the absolute is greater than or equal to zero we leave it alone, if the value inside the value is less than zero we take the negative of it

OpenStudy (anonymous):

so for x-(1/5)>0 the absolute value can just be removed from the equation. x-(1/5)=3 can you tell me what the equation is for x-(1/5)<0 is?

OpenStudy (anonymous):

oo i think i got the answer

OpenStudy (anonymous):

what do you think the answer is?

OpenStudy (anonymous):

d. x = 14/5 and x = 16/5

OpenStudy (anonymous):

you may want to double check 14/5 I think you made a very tiny error.

OpenStudy (anonymous):

oh so 14/5 would be a negative than

OpenStudy (anonymous):

yes now that your question is answered please close this one and post your next question as a separate question

OpenStudy (anonymous):

also medals are always appreciated

OpenStudy (anonymous):

Expand the LHS (Left Hand Side). \[3x-10=3x-12+2\] \[3x-10=3x-10\] Therefore if you put any number for x you get the same thing. That means this equation holds true for every single x value. Thus x=infinitely many solutions.

OpenStudy (anonymous):

actually @DrPepperx3 , @Azteck is right I didn't look at the question the other person helped you with closely until now